OpenStudy (anonymous):

convert fraction 0.125

OpenStudy (campbell_st):

start with 125/1000 and simplify

OpenStudy (anonymous):

yes i dont know how to simpluify

Directrix (directrix):

125 = 5 x 5 x 5 where x is multiply (not a variable) 1000 = 2 x 2 x2 x 5 x 5 x 5

OpenStudy (anonymous):

0.125 simplyfy to lowest term fraction

Directrix (directrix):

So, the next procedure is to divide out like factors. 5 x 5 x 5 all over 2 x 2 x2 x 5 x 5 x 5 = 1 /( 2 x 2 x2) = 1 / 8

OpenStudy (anonymous):

omg i need a math tutor

OpenStudy (anonymous):

ty

Directrix (directrix):

You're okay. Probably have a case of the heebie-jeebies. It will pass.

Directrix (directrix):

Look and see if the answer makes sense. If not, I can draw it out. BTW, if you are in college or some other learning place, there's most likely a tutoring service that is free. That is, face-to-face.

OpenStudy (anonymous):

yea havnt been in school in 20 years

Directrix (directrix):

At least you're back. That takes guts.

OpenStudy (anonymous):

fraction ratio of 8 to 9 lowest terms

Directrix (directrix):

8/9. 8 and 9 have no common factors other than 1. There's nothing to divide out. 8/9 is already in lowest terms.